Circular No. 8313

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 F2 (LINEAR) The LINEAR project reports the discovery of a comet with a tail in p.a.

≈ 235^{o} on their images from Mar. 26.5 UT (discovery observation below). Following posting on the NEO Confirmation Page, P. Holvorcem reports that three co-added unfiltered CCD images taken with the Tenagra 0.81-m telescope on Mar. 29.48 show the object to be diffuse with a 10′′ coma.

2004 UT  α_{2000}  δ_{2000}  Mag.
Mar. 25.43956  18^{h}26^{m}21.^{s}07  −4^{o}24′10′′.0  18.4

The available astrometry, the following parabolic orbital elements, and an ephemeris appear on MPEC 2004-F81.

T = 2003 Dec. 26.945 TT  ω = 317.^{o}249
Ω = 248.280 }2000.0
q = 1.43092 AU  i = 104.879

COMET P/2004 F3 (NEAT) The NEAT program reports the discovery of a comet on images

taken with the 1.2-m reflector at Haleakala (discovery observation below), with a tail ∼ 5′′ long toward the west. Following WWW posting on the NEO Confirmation Page, numerous observers have reported on the object’s cometary appearance on CCD images taken during Mar. 28.9–29.4 UT, including F. Hormuth, J. Tichá and M. Tichý, P. Kušnirak, B. L. Stevens, J. Young, and G. Hug — the object generally showing a coma diameter of 8′′–30′′, total magnitude as bright as 16, and a faint tail ≈ 20′′–40′′ long spanning p.a. 240^{o}–285^{o}.

2004 UT  α_{2000}  δ_{2000}  Mag.
Mar. 28.54336  13^{h}43^{m}27.^{s}00  +11^{o}14′04′′.3  17.1

The available astrometry, the following elliptical orbital elements, and an ephemeris appear on MPEC 2004-F82. The comet passed ∼ 0.37 AU from Jupiter in July 2001, causing the perihelion distance to decrease.

T = 2005 Jan. 7.465 TT  ω = 176.^{o}433
e = 0.28517  Ω = 78.867 }2000.0
q = 2.87248 AU  i = 16.009
a = 4.01840 AU n^{o} = 0.122356 P = 8.1 years
2004 March 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green